1. Negative O 2. Positive O 3. Unable to determine Submit answer Question (part 1 of 2) A small 1.02 g plastic ball is suspended by a 21.8 cm long string at an angle of 10.7 in a uniform electric field of 1530 N/C, as shown. 10.7 1530 N/C \21,8 cm 1.02 g Is the balls charge positive or negative? Question (part 2 of 2) Submit answer Your response... If the ball is in equilibrium when the string makes a 10.7 angle with the vertical, what is the net charge on the ball? The acceleration due to gravity is 9.8 m/s, and the Coulomb constant is 8.98755 x10 N m2/C2 Answer in units of fc.
Answer #1

given

mass m = 1.02 g

= 1.02 x 10-3 kg

length L = 21.8 cm

= 0.218 m

angle \theta = 10.7o

E = 1530 N/C

g = 9.8 m/sec2  

q = ?

using equation F = m g tan\theta

and F = q E

equating both equations

m g tan\theta = q E

q = m g tan\theta / E

= 1.02 x 10-3 x 9.8 x tan10.7 / 1530

q = 0.00000123449 C

or

q = 1.23449 x 10-6\muC